Suppose you were given Php 1,000,000,000. How long would it take to have less than five pesos left if you spent half of it every day?

1
Expert's answer
2021-11-17T05:06:01-0500

Assuming that he is spending half of what is left everyday


After first day reminder is "=\frac{1}{2}×10^9"


"n^{th}" day the reminder will be "=(\frac{1}{2})^{n}×10^9<5"


"\frac{10^{9}}{5}<(\frac{1}{2})^{-n}"


Taking logs both sides


log"(2×10^{8})<-n\>log\>0.5"


"8.3010<0.3010n"

"27.58<n"


"\therefore" Number of days "=28 days"